# Pushing the Limits of Distillation-Based Class-Incremental Learning via Lightweight Plugins

### v2

@@ Abstract
Existing replay and distillation-based class-incremental learning (CIL) methods are effective at retaining past knowledge but are still constrained by the stability–plasticity dilemma. Since their resulting models are learned over a sequence of incremental tasks, they encode rich representations and can be regarded as pre-trained bases. Building on this view, we propose a plug-in extension paradigm termed Deployment of LoRA Components (DLC) to enhance them. For each task, we use Low-Rank Adaptation (LoRA) to inject task-specific residuals into the base model’s deep layers. During inference, representations with task-specific residuals are aggregated to produce classification predictions. To mitigate interference from non-target LoRA plugins, we introduce a lightweight weighting unit. This unit learns to assign importance scores to different LoRA-tuned representations. Like downloadable content in software, DLC serves as a plug-and-play enhancement that efficiently extends the base methods. Remarkably, on the large-scale ImageNet‑100, with merely 4% of the parameters of a standard ResNet-18, our DLC model achieves a significant 8% improvement in accuracy, demonstrating exceptional efficiency. Under a fixed memory budget, methods equipped with DLC surpass state-of-the-art expansion-based methods.
@@ 1 Introduction
In recent years, deep learning has achieved remarkable success across a variety of domains [14, 36, 5, 4, 34]. However, most models rely on a key assumption that the data distribution and learned classes remain fixed. In real-world scenarios, data distributions evolve continuously and new classes emerge over time [11], making static training paradigms impractical. Moreover, learning each new task typically relies on additional task-specific sub-models, leading to inefficient model growth over time. Class-Incremental Learning (CIL) [42, 17, 35] thus plays a crucial role in enabling models to continually learn from non-stationary data streams while retaining prior knowledge.
@@
Existing CIL methods  [42] can be broadly categorized into rehearsal-based, restriction-based, and expansion-based approaches. Rehearsal-based methods maintain a small replay buffer by storing a limited number of representative exemplars from previous tasks [18, 24, 8]. Other categories of methods likewise adopt rehearsal mechanisms to retain exemplars from past tasks, thereby further improving model accuracy. Expansion-based CIL methods [19, 25, 32] allocate task-specific architectural blocks, creating isolated parameter spaces for new knowledge while freezing previously learned parameters. This strategy alleviates catastrophic forgetting and preserves plasticity on new tasks. However, rather than updating a single model incrementally, these methods expand the network and store, per task, an additional full [33] or partial (about 70%) [43] feature extractor, incurring substantial storage costs. For instance, with ResNet-18, each new task adds about 11.7MB of parameters, roughly equivalent to storing 3,998 images, which is about twice the replay buffer capacity under standard CIL settings. This overhead can limit their practicality for long task sequences, resource-constrained devices, and large-scale deployments.
@@
Figure 1: Parameter-accuracy comparison of different CIL methods on ImageNet-100 with ResNet-18. DLC achieves the best balance between parameter efficiency and accuracy compared to prior SOTA methods. iCaRL-DLC, WA-DLC, and BiC-DLC mean integrating DLC into the corresponding methods, respectively.
@@
Restriction-based methods [21, 2] aim to mitigate catastrophic forgetting by constraining model updates on the new task. A representative example is knowledge distillation [23, 26], which preserves previous knowledge by aligning the current model’s output with its earlier versions. These methods perform CIL on a single model, leading to substantially lower storage overhead. As a result, these approaches are practically attractive in real-world settings. However, they inevitably introduce a trade-off [7]: the model must balance adherence to past knowledge with adaptation to new data. This stability-plasticity dilemma often becomes the key bottleneck for distillation-based approaches. Many enhancements [31, 38, 10] have been proposed to improve their accuracy, yet the core issue persists: distillation and new-task fitting are coupled objectives whose gradients can conflict, leaving performance governed by the same trade-off. A natural way to reduce this interference is to expand extra parameters for new tasks. However, scaling up the feature extractor is quickly parameter-prohibitive, and typical expansion-based schemes retrain task-specific extractors, breaking the shared single-model predictor that distillation relies on. This exposes a new quandary to CIL: How can we break the stability–plasticity dilemma by introducing a small set of task-specific auxiliary parameters, while remaining friendly to distillation-based methods?
@@
To address this challenge, we propose the Deployment of LoRA Components (DLC), a plug-and-play extension paradigm for CIL. We revisit parameter-efficient tuning modules in transfer learning beyond the frozen feature extractor assumption, and study them under the training dynamics of replay and distillation-based CIL. We attach task-specific lightweight plugins to deep layers of the base model and instantiate each plugin with LoRA in this work. These plugins provide additional knowledge by injecting residuals into the extracted features. For each incoming task, we first follow the underlying baseline to train the backbone while keeping all plugins frozen, and then update only the plugin associated with the current task. This decoupled update strategy trains each plugin independently, enabling DLC to be deployed into the baseline without interference. During inference, the model activates all plugins to extract their feature representations, which are aggregated for final classification. However, plugins that are not aligned with the input sample may contribute noisy or weakly informative residuals. We introduce a lightweight weighting unit before the classifier, which automatically assigns importance weights to different plugin outputs. Experimental results on multiple benchmarks show that DLC consistently improves replay and distillation-based baselines. For example, DLC achieves up to 8% accuracy gains on ImageNet-100 with ResNet-18 and on CIFAR-100 with ResNet-32, while increasing the total parameter count by only 4% to 30%. Under the same memory budget, DLC also outperforms existing state-of-the-art CIL baselines. Our contributions are summarized as follows:
@@
  • We propose DLC, a plug-and-play plugin extension framework that enables lightweight task-specific plugin residual injection in continually updated backbones, anchored by the consistent knowledge preserved through replay and distillation.
@@
  • We introduce a lightweight weighting unit that adaptively suppresses irrelevant plugin activations based on learned importance, reducing feature interference across tasks and leading to more stable and accurate incremental classification.
@@
  • Extensive experiments show that DLC significantly boosts various CIL methods’ performance. Meanwhile, under the fixed memory budget, the DLC-enhanced distillation method could surpass SOTA CIL methods.
@@ 2 Related Work
2.1 Class Incremental Learning
Class-incremental learning primarily aims to prevent the model from forgetting old classes. Existing methods tackle this challenge from multiple perspectives, including rehearsal [16, 1] that replays representative exemplars from previous classes, regularization [3] or distillation [12] that limit overly aggressive updates on new data, expansion-based designs that train and retain task-specific full [33] or partial [43] feature extractors, and further integrate contrastive objectives [39] to stabilize representations. In addition, some methods leverage transfer learning with Transformers pretrained on large-scale datasets to provide stronger representations [40, 27]. These methods often exhibit an accuracy–memory trade-off: approaches with higher recognition accuracy typically require storing more replay samples or allocating and maintaining additional model parameters through expansion or larger backbones.
2.2 Distillation-Based CIL
Distillation-based CIL trains the model on new tasks while regularizing it to match the predictions of the previous version. A wide range of distillation-based approaches have been explored, including combining distillation with replay [23], incorporating bias correction [31] or weight alignment [38], momentum-based distillation [20], and plug-and-play auxiliary objectives [10]. Distillation remains appealing as it introduces minimal overhead, can leverage unlabeled incoming samples, and fits a simple single-model training loop, making it broadly applicable in practical settings such as online [30], imbalanced [22], and few-shot [37] CIL. However, the feature extractor’s fixed capacity imposes a fundamental constraint: the model must balance stability and plasticity by trading off classification and distillation losses, ultimately capping achievable accuracy.
Figure 2: The proposed DLC framework. Left: Training. When a new task $t$ arrives, a dedicated plugin set $L_{t}$ is created. Then sequentially train the feature extractor and $L_{t}$. Right: Test. The feature extractor sequentially loads all task plugins $L_{t}$ to produce enhanced representations, which are concatenated, weighted by a gating unit, and classified.
@@ 3 Preliminaries
3.1 Problem Formulation of CIL
A CIL model aims to learn new classes while retaining previous knowledge from the continuously evolving data streams $D_{1},D_{2},\cdots,D_{t}$ [7], where $D_{i}=\{(\mathbf{x}_{j},y_{j})\}_{j=1}^{n_{i}}$ containing $n_{i}$ instances. One class will appear only once on a given $D_{i}$ and will not be repeated. At each incremental learning stage $t$, only the current data stream $D_{t}$ and a few exemplars from the previous classes [23] can be used for model training. For any stage $t$, the model has learned $D=D_{1}\cup D_{2}\cup\cdots\cup D_{t}$ streams, We aim to find a model $f(\mathbf{x}):\mathcal{X}\rightarrow\mathcal{Y}$ that minimizes empirical risk across all learned classes, where $\mathcal{X}$ and $\mathcal{Y}=Y_{1}\cup Y_{2}\cdots\cup Y_{t}$ represent the sample’s feature and class spaces, respectively. The effectiveness of the model is evaluated by the test dataset $D^{test}=D_{1}^{test}\cup D_{2}^{test}\cup\cdots\cup D_{t}^{test}$, which can be expressed in Eq.(1):
@@ 3.1 Problem Formulation of CIL
Consistent Knowledge Preserved in CIL Considering the classic CIL scenario with only a single feature extractor. The objective function used in non-CIL settings is the cross-entropy loss $\mathcal{L}_{\text{CE}}$, while restriction-based CIL methods incorporate an additional restriction term $\mathcal{L}_{\text{MEM}}$ that enforces the preservation of knowledge acquired from previous tasks. This composite objective function can be formally expressed as:
$\mathcal{L}=\mathcal{L}_{\text{CE}}+\mathcal{L}_{\text{MEM}}.$ (2)
As a canonical example of restriction-based methods, we examine logit-based distillation, which ensures consistency between the logit distributions produced by the old and new models for known classes. Specifically, the restriction term $\mathcal{L}_{\text{MEM}}$ is defined as a distillation loss $\mathcal{L}_{\text{KD}}$. Let $\hat{q}(\boldsymbol{x})$ and $q(\boldsymbol{x})$ represent the logits of the teacher and student models. The cross-entropy version of $\mathcal{L}_{\text{KD}}$ is defined as:
@@
where the $\tau$ is the temperature scalar.
Moreover, most methods are complemented by a replay buffer that stores a subset of representative exemplars from previous tasks. These exemplars are interleaved with data from the current task during training to help maintain performance on past tasks. Specifically, when learning the $t$-th task, the model is conventionally trained only on the new dataset $D_{t}$. With the replay buffer, the training dataset is supplemented with $D_{\text{exp}}$ retrieved from the buffer, forming a combined training set $D_{\text{train}}=D_{t}\cup D_{\text{exp}}$.
Although the model’s parameters are dynamically updated throughout the incremental learning process, the use of knowledge distillation and the replay buffer ensures that a portion of knowledge from previous tasks remains preserved. Replay ensures the model continues to minimize the empirical risk on a representative subset of past tasks, anchoring the parameters in regions consistent with previous data. Concurrently, knowledge distillation functions as a functional regularizer, ensuring that the mappings for old classes remain stable by aligning the output logits (or distributions) with those of the prior model. Consequently, a model trained with replay and distillation preserves substantial knowledge of past classes, resulting in a single unified model that is broadly effective but suboptimal. We posit that the knowledge stored in the model can be efficiently activated with minor task-specific parameters.
@@ 4 The Proposed Method
DLC proposes a plugin-level residual injection paradigm, as illustrated in Fig. 2. We treat a replay and distillation-trained network as a base model that accumulates knowledge across tasks but remains suboptimal in accuracy, and then augment it with task-specific plugins that inject residual knowledge into the extracted features to improve performance. In this work, each plugin is instantiated with a LoRA-style low-rank parameterization, but is used as a lightweight task extension rather than a conventional fine-tuning update. The proposed DLC is detailed in the following subsections.
4.1 Task-Specific Plugins Deployment
Parameter-efficient tuning is typically used in transfer learning by adapting a feature extractor with frozen parameters, where layer inputs remain largely stationary. Here, we revisit it beyond this static-backbone assumption and study it under replay- and distillation-based class-incremental learning, where old-task outputs are constrained, and intermediate feature drift is therefore controlled during continual updates. Concretely, after training on task $t$, the output deviation at layer $\ell$ is bounded by $K_{\ell}\Gamma_{t}$ (see Appendix), where $K_{\ell}$ depends on hyperparameters such as network structure and distillation temperature, and $\Gamma_{t}$ reflects the distribution discrepancy induced by the chosen replay and distillation strategy. This bound directly limits the worst-case feature drift on the old-task distribution, enabling stable optimization of lightweight auxiliary parameters in this dynamic regime.
We introduce a plugin extension design and maintain an independent plugin set for each task. Specifically, for each task $t$, we create a dedicated set of $k$ plugins for $\phi(\boldsymbol{x})$, denoted as $L_{t}$. The $\phi(\boldsymbol{x})$ equipped with $L_{t}$ is denoted by $\phi^{L_{t}}(\boldsymbol{x})$. In DLC, plugins are deployed in the final layers of $\phi(\boldsymbol{x})$. Taking a CNN-based architecture as an example, the first plugin in $L_{t}$ is attached to the last convolutional layer of $\phi(\boldsymbol{x})$, and the remaining plugins are placed on progressively earlier layers moving backward through the network. This design is motivated by two considerations. First, deeper features are more semantically meaningful and are directly used for classification, so injecting task-specific residuals at this level can more effectively enhance inter-class discriminability. Second, it is more compatible with distillation-based objectives. If plugins are deployed in shallow layers, residual changes in low-level features would propagate through the network and amplify logit deviations, increasing the optimization difficulty of distillation losses that rely on output alignment between old and new models.

### v3

# Pushing the Limits of Distillation-Based Continual Learning via Classifier-Proximal Lightweight Plugins

@@ Abstract
Continual learning requires models to learn continuously while preserving prior knowledge under evolving data streams. Distillation-based methods are appealing for retaining past knowledge in a shared single-model framework with low storage overhead. However, they remain constrained by the stability-plasticity dilemma: knowledge acquisition and preservation are still optimized through coupled objectives, and existing enhancement methods do not alter this underlying bottleneck. To address this issue, we propose a plugin extension paradigm termed Distillation-aware Lightweight Components (DLC) for distillation-based CL. DLC deploys lightweight residual plugins into the base feature extractor’s classifier-proximal layer, enabling semantic-level residual correction for better classification accuracy while minimizing disruption to the overall feature extraction process. During inference, plugin-enhanced representations are aggregated to produce classification predictions. To mitigate interference from non-target plugins, we further introduce a lightweight weighting unit that learns to assign importance scores to different plugin-enhanced representations. DLC could deliver a significant 8% accuracy gain on large-scale benchmarks while introducing only a 4% increase in backbone parameters, highlighting its exceptional efficiency. Moreover, DLC is compatible with other plug-and-play CL enhancements and delivers additional gains when combined with them.
@@ 1 Introduction
In recent years, deep learning has achieved remarkable success across a broad spectrum of multimedia applications [39, 4, 10, 35, 41]. However, most existing models are built on the assumption that data distributions and semantic categories remain fixed after training. In real-world multimedia scenarios, data streams evolve continuously, and novel categories emerge over time [9], making static training paradigms increasingly inadequate. Meanwhile, naively adapting models to newly arrived data often requires additional specific modules, leading to inefficient model growth and limited scalability. Continual learning (CL) [47, 20, 40] therefore plays a crucial role in enabling deep learning models to continuously adapt to non-stationary data while preserving prior knowledge.
@@
Existing CL methods [47] can be broadly categorized into rehearsal-based, restriction-based, and expansion-based approaches. Rehearsal-based methods maintain a small replay buffer by storing a limited number of representative exemplars from earlier learning stages [22, 28, 6]. Other categories of methods likewise adopt rehearsal mechanisms to retain exemplars from previous learning stages, thereby further improving model accuracy. Expansion-based CL methods [23, 29, 37] allocate stage-specific architectural blocks, creating isolated parameter spaces for newly acquired knowledge while freezing previously learned parameters. This strategy alleviates catastrophic forgetting and preserves plasticity during subsequent learning stages. However, rather than updating a single model progressively, these methods expand the network and store, for each learning stage, an additional full [38] or partial (about 70%) [48] feature extractor, incurring substantial storage costs. For instance, with ResNet-18, each new learning stage adds about 11.7MB of parameters, roughly equivalent to storing 3,998 images, which is about twice the replay buffer capacity under standard CL settings. This overhead can limit their practicality for continual learning.
@@
Figure 1: Parameter-accuracy comparison of different CL methods on ImageNet-100. DLC achieves the best balance between parameter efficiency and accuracy compared to prior distillation-based and SOTA expansion-based methods. iCaRL-DLC, WA-DLC, and BiC-DLC mean integrating DLC into the corresponding methods, respectively.
@@
Restriction-based methods [25, 2] aim to mitigate catastrophic forgetting by constraining model updates during new learning stages. A representative example is knowledge distillation [27, 30], which preserves previous knowledge by aligning the current model’s output with its earlier versions. These methods perform CL on a single model, leading to substantially lower storage overhead. As a result, these approaches are practically attractive in real-world settings. However, they inevitably introduce a trade-off [5]: the model must balance adherence to past knowledge with adaptation to new data, since the distillation and knowledge acquisition are coupled objectives whose gradients can conflict. This stability-plasticity dilemma often becomes the key bottleneck for distillation-based approaches. Many enhancement methods [36, 43, 8] have been proposed to improve their accuracy, but most of them mainly introduce additional training objectives, leaving the core stability-plasticity dilemma fundamentally unresolved, leaving performance governed by the same trade-off. A natural way to reduce this interference is to expand extra parameters. However, scaling up the feature extractor is quickly parameter-prohibitive, and expanding multiple extractors breaks the shared single-model predictor that distillation relies on. This exposes a new quandary to CL: How can we break the stability–plasticity dilemma by introducing a small set of auxiliary parameters, while remaining friendly to distillation-based methods?
@@
To address this challenge, we propose the Distillation-aware Lightweight Components (DLC), a plug-and-play extension paradigm for CL. We revisit parameter-efficient tuning modules in transfer learning and study them under the training dynamics of distillation-based CL, where the backbone is not a knowledge-rich pretrained model but a dynamically evolved learner trained over continual stages. We attach several lightweight plugins to the classifier-proximal layer of the base feature extractor, i.e., its final feature layer before classification. These plugins perform specific semantic calibration by injecting residual corrections into classifier-proximal features. When adding each plugin, we first follow the underlying baseline to train the backbone while keeping all plugins frozen, and then freeze the backbone to train only the plugin introduced for the current stage. This decoupled update strategy trains each plugin independently, enabling DLC to be deployed into the baseline without interference. During inference, the model activates all plugins to extract their feature representations, which are aggregated for final classification. However, plugins that are not aligned with the input sample may contribute noisy or weakly informative residuals. We introduce a lightweight weighting unit before the classifier, which automatically assigns importance weights to different plugin outputs. Experimental results on multiple benchmarks show that DLC consistently improves replay and distillation-based baselines. For example, DLC achieves up to 8% accuracy gains on ImageNet-100 with ResNet-18 and on CIFAR-100 with ResNet-32, while increasing the total parameter count by only 4% to 30%. Under the same memory budget, DLC also outperforms existing state-of-the-art CL baselines.
Our contributions are summarized as follows:
@@
  • We propose DLC, a plug-and-play plugin extension framework that deploys lightweight residual plugins at classifier-proximal layers in continually updated backbones, enabling effective semantic correction while preserving compatibility with distillation-preserved knowledge.
@@
  • We introduce a lightweight weighting unit that adaptively suppresses irrelevant plugin activations based on learned importance, leading to more stable and accurate continual learning.
@@
  • Extensive experiments show that DLC significantly boosts the performance of various CL methods and remains compatible with other plug-and-play enhancement techniques. Meanwhile, under a fixed memory budget, the DLC-enhanced distillation method can surpass state-of-the-art expansion-based CL methods.
@@ 2 Related Work
2.1 Continual learning
Continual learning primarily aims to prevent the model from forgetting acquired knowledge. Existing methods tackle this challenge from multiple perspectives. Rehearsal-based methods [19, 1] replays representative exemplars from previous classes. Regularization [3] or distillation-based methods [27, 36, 43, 13] limit overly aggressive updates on new data. Expansion-based methods train and retain task-specific full [38] or partial [48] feature extractors, and further integrate contrastive objectives [44] or pre-trained models [45, 31] for better representations. These methods often exhibit a performance–memory trade-off: approaches with higher overall performance typically require storing more replay samples or allocating and maintaining additional model parameters through expansion.
2.2 Distillation-Based CL
Distillation-based CL trains the model on new data while regularizing it to match the predictions of the previous version. A wide range of distillation-based approaches have been explored, including combining distillation with replay [27], incorporating bias correction [36] or weight alignment [43], momentum-based distillation [24], and plug-and-play enhancement methods [3, 8, 18]. Distillation remains appealing as it introduces minimal overhead, can leverage unlabeled incoming samples, and fits a simple single-model training loop, making it broadly applicable in practical settings [34, 26, 42]. However, the feature extractor’s fixed capacity imposes a fundamental constraint: the model must balance stability and plasticity by trading off new-data learning and distillation losses, ultimately capping achievable performance.
Figure 2: The proposed DLC framework. Left: Training. When a new $D_{t}$ arrives, a dedicated plugin $L_{t}$ is created. First follows the base method to train the backbone, and then freezes the backbone to train $L_{t}$. Right: Test. Using all plugins to enhance the feature extractor representations, then reweight by a weighting unit, concatenated, and classified.
@@ 3 Preliminaries
3.1 Problem Formulation
A CL model aims to learn new data while retaining previous knowledge from the continuously evolving data streams $D_{1},D_{2},\cdots,D_{t}$, where $D_{i}=\{(\mathbf{x}_{j},y_{j})\}_{j=1}^{n_{i}}$ containing $n_{i}$ instances. At each learning stage $t$, the current data stream $D_{t}$ is treated as a task, and training is performed using only $D_{t}$ with a few exemplars from the previous stage [27]. After the model advances to the next learning stage, all earlier stages, except for the stored exemplars, will become inaccessible. When there are test samples for inference, the model needs to make predictions based on the label spaces of all the tasks it has encountered. We do not provide the task ID of the sample to the model, making this a more challenging class-incremental continual learning [5] setup.
For any stage $t$, the model has learned $D=D_{1}\cup D_{2}\cup\cdots\cup D_{t}$ tasks, We aim to find a model $f(\mathbf{x}):\mathcal{X}\rightarrow\mathcal{Y}$ that minimizes empirical risk across all learned data, where $\mathcal{X}$ and $\mathcal{Y}=Y_{1}\cup Y_{2}\cdots\cup Y_{t}$ represent the sample’s feature and label spaces, respectively. The effectiveness of the model is evaluated by the test dataset $D^{test}=D_{1}^{test}\cup D_{2}^{test}\cup\cdots\cup D_{t}^{test}$, which can be expressed in Eq.(1):
@@ 3.1 Problem Formulation
Knowledge in Distillation-Based CL Considering the classic CL scenario with only a single feature extractor. The objective function used in non-CL settings is the cross-entropy loss $\mathcal{L}_{\text{CE}}$, while distillation-based CL methods incorporate an additional restriction term $\mathcal{L}_{\text{KD}}$ that enforces the preservation of knowledge acquired from previous data. This composite objective function can be formally expressed as:
$\mathcal{L}=\mathcal{L}_{\text{CE}}+\mathcal{L}_{\text{KD}}.$ (2)
As a canonical example of distillation-based methods, we examine logit-based distillation, which ensures consistency between the logit distributions produced by the old and new models for known classes. Specifically, let $\hat{q}(\boldsymbol{x})$ and $q(\boldsymbol{x})$ represent the logits of the teacher and student models. The cross-entropy version of $\mathcal{L}_{\text{KD}}$ is defined as:
@@
where the $\tau$ is the temperature scalar. Most of these methods are complemented by a replay buffer that stores a subset of representative exemplars from previous data. These exemplars are interleaved with data from the current stage during training to help maintain performance on past data. Specifically, in the conventional setting, the model is trained only on the newly arrived data $D_{t}$. With the replay buffer, the training data is supplemented with $D_{\text{exp}}$ retrieved from the buffer, forming a combined training set $D_{\text{train}}=D_{t}\cup D_{\text{exp}}$. Although the model’s parameters are dynamically updated throughout the continual learning process, the use of knowledge distillation and the replay buffer ensures that a portion of knowledge from previous data remains preserved. Replay ensures the model continues to minimize the empirical risk on a representative subset of past data, anchoring the parameters in regions consistent with previous data. Concurrently, knowledge distillation functions as a functional regularizer, ensuring that the mappings for old classes remain stable by aligning the output logits (or distributions) with those of the prior model. Consequently, a model trained with replay and distillation preserves substantial knowledge of past classes, resulting in a single unified model that is broadly effective but suboptimal. We posit that the knowledge stored in the model can be efficiently activated with minor auxiliary parameters.
@@ 4 The Proposed Method
DLC proposes a plugin-level residual injection paradigm, as illustrated in Fig. 2. We treat a distillation-based CL network as the base model that accumulates knowledge across each CL stage but remains suboptimal in accuracy, and then augment it with auxiliary plugins that inject residual knowledge into the extracted features to improve performance. The proposed DLC is detailed in the following subsections.
4.1 Classifier-Proximal Plugin Deployment
We introduce a plugin extension design for continual learning by maintaining an independent residual plugin for each learning stage. Specifically, for stage $t$, we attach a dedicated plugin $L_{t}$ for the feature extractor $\phi(\boldsymbol{x})$, and denote the resulting plugin-enhanced representation by $\phi^{L_{t}}(\boldsymbol{x})$. Rather than modifying the backbone weights directly, each plugin acts as a lightweight residual branch that performs stage-specific semantic correction on the extracted features. In this way, DLC complements the continually updated backbone with additional discriminative knowledge in a plug-and-play manner, while preserving the backbone as the main carrier of shared knowledge across stages.
DLC adopts a classifier-proximal deployment strategy, where plugins are attached to the final layer of the $\phi(\boldsymbol{x})$, i.e., the layer closest to the classifier. This placement is beneficial for two reasons. First, classifier-proximal features are already semantically mature and directly support classification, so residual correction at this stage can more effectively improve inter-class discriminability. Second, the injected perturbation is localized near the classifier and thus minimally disrupts the overall feature extraction process. Compared with shallow-layer intervention, such a design is more compatible with distillation-based objectives, since shallow-layer feature changes would otherwise propagate through the network and amplify output deviations, making output alignment between the old and current models harder to optimize.